OpenStudy (anonymous):

given 2 angles and a side (80 degrees,20 degrees, and 8 cm) how do you find the other two sides (round 3 decimal places)

OpenStudy (anonymous):

hmm I'm not sure , but I believe that the following question represents a graph with 3 dimenstions where: x = r cos (theta) y = r sin (theta) z = r? or r cos (theta) <-- not sure of this one. If z = r = 8 cm, then you can plug in the R in the x and y equations to find the other sides; however, if r is not = z, then I'm not sure of what to do. Hope I'm right though :)

OpenStudy (dumbcow):

is this for a triangle?

OpenStudy (dumbcow):

if so use law of sines sinA/a = sinB/b = sinC/c where side a is opposite angle A

OpenStudy (dumbcow):

cant solve this for you because i dont know where the angles and sides are relative to each other you find the 3rd angle by subtracting from 180 180 - 80-20 = 80 3rd angle = 80 degrees looks like an isosceles triangle

OpenStudy (anonymous):

>_< I've made it so complicated lol, thank you dc ^_^